# Incommensurate and commensurate antiferromagnetic orders in the kagome compound UV6Sn6

### v1

@@ III.2 Magnetization
The magnetization data collected from a single crystal of UV6Sn6, with a field of 1 T applied parallel and perpendicular to the crystallographic $c$-axis, is presented in Fig. 3. The magnetization as a function of temperature reveals a strong magnetic anisotropy, with the $c$-axis being the easy magnetization axis (Fig. 3a). Along the $ab$-plane, the magnetization remains low and nearly constant across the entire temperature range. Notably, the magnetization parallel to the $c$-axis exhibits a clear maximum at $T_{N_{1}}\approx 30$ K. A second transition occurring at approximately 23 K is also revealed as a kink in the magnetization curve (inset Fig. 3a). The effective magnetic moment, derived from the Curie-Weiss fit, which holds down to about 70 K, (shown in Fig. 3b), is found to be $3.89(1)$$\mu_{B}$/U. This value is slightly higher than the effective moment of the U3+ or the U4+ configurations. From the positive Curie-Weiss temperature $\Theta_{CW}=27.8$ K a dominant ferromagnetic exchange at high temperatures is expected.
The magnetization isotherms $M(H)$ collected at 2 K are shown in Fig. 3c. When the magnetic field is applied parallel to the $c$-axis, a clear hysteresis effect is observed, along with two notable changes in the magnetization. There is a sharp change in slope at approximately 3 T, followed by a tendency towards saturation starting around 4 T. The magnetization at 6 T is roughly 1 $\mu_{B}$, which is lower than the ordered moment of the free U ion. However, as discussed below, the magnetization at 6 T may not correspond to the fully polarized moment.
Figure 3: (a) Temperature dependence of the magnetization under an applied magnetic field of 1 T parallel and perpendicular to the $c$-axis. Inset: enlarged view at low temperatures. The arrows indicate the anomalies at the two transitions. (b) Temperature dependence of the inverse susceptibility at 1 T along the c-axis. The dashed line is a fit to the Curie-Weiss law. (c) Magnetic field dependence of the magnetization at 2 K for the field parallel and perpendicular to the $c$-axis.
@@ III.3 Resistivity
The temperature dependence of the normalized electric resistivity at magnetic fields of 0 and 9 T applied along the $c$-axis is presented in Fig. 4. The resistivity ratio $RRR=R(300{\rm K})/R(1.8{\rm\,K})\approx 3$ is rather low. On cooling the resistivity first increases slightly and shows a maximum around 180 K. At high temperatures, no magnetoresistance occurs and down to 90 K the temperature dependence at 9 T is similar to that at zero magnetic field. On further cooling, the resistivity shows a marked kink at the onset of the magnetic order near $T_{N1}\simeq 29~{}K$ (see Fig. 4b). At 6 T the magnetic transition is still visible by a clear kink, while it disappears at 9 T.
Figure 4(c) presents the derivative of the normalized resistivity as a function of temperature. It shows a mean-field like anomaly at the transition as expected from the Fisher-Langer behavior $d\rho/dT\propto C_{P}$, often observed in metallic systems [32, 33]. In particular, we do not observe an increase of resistivity upon lowering the temperature through the magnetic transition, unlike other incommensurate antiferromagnets due to the opening of a superzone gap in the electronic band structure [34, 35, 36, 37, 38, 39]. This may be due to the electrical current direction being perpendicular to the magnetic wave-vector (see neutron scattering and the magnetic structure discussion below).
@@
Figure 4: (a) Temperature dependence of the electrical resistivity at magnetic fields of 0 T and at 9 T applied along the $c$-axis, normalized to the value at 300 K. (b) Zoom of the low temperature range for fields of 0, 6, and 9 T. The arrows indicate the magnetic ordering temperature $T_{N1}$. (c) Temperature derivative of the resistivity at zero field as a function of temperature. The arrows shows to the magnetic transition temperatures $T_{N1}$ and $T_{N2}$.
@@ III.4 Heat Capacity
Figure 5: Specific heat divided by temperature vs temperature of UV6Sn6 at zero magnetic field. The arrows indicate the position of $T_{N1}=28.8$ K and $T_{N2}=23$ K. The inset shows $C_{p}$ as a function of $T^{2}$. The dashed line is a linear fit to $C_{p}/T=\gamma+\beta T^{2}$ and we determine $\gamma=15$mJmol-1K-2 and $\beta=2.95$ mJmol-1K-4.
Figure 6: (a) Specific heat divided by temperature at selected T values as a function of increasing magnetic fields. (b) Field dependence of the specific heat at 10 K in an enlarged view. (c) Magnetic phase diagram for field $H\parallel c$ from the field dependent specific heat measurements (increasing fields). (Lines are guides to the eyes.)
@@
In Fig. 6(a) we show the field dependence of the specific heat at various temperatures for increasing magnetic fields. As shown explicitly in Fig. 6(b) for the curve at 10 K, a first increase of the specific heat is observed at 2.5 K while a very marked anomaly occurs at 7.7 K. While the low field anomaly for all temperatures is very small, the high field anomaly appears as a very pronounced peak with a first-order like behavior at 15 and 20 K. From these anomalies, a magnetic phase diagram can be proposed as illustrated in Fig. 6(c). The low field anomaly is associated with a change in the magnetic structure, as discussed below.
@@ III.5 Neutron Scattering and Magnetic Structure
Mapping of the reciprocal space along the high-symmetry lines of the Brillouin zone was carried out on D23 with the copper monochromator in order to identify the magnetic ordering wave-vector at 1.41 K. The outcome is the identification of a single wave-vector that is described below. Given the complexity of the UV6Sn6 crystal structure (Fig. 1), we approach the magnetic structure characterization from the HfFe6Ge6-type parent cell. Figure 7 shows scans around the momentum transfer ${\bf Q}=(1,0,0.5)$ along the directions $Q_{H}$ (panel a), $Q_{K}$ (panel b) and $Q_{L}$ (panel c) respectively, at 1.5 K and 30 K. The coordinates of the momentum transfer $\bf{Q}$ are given in reciprocal lattice units (r.l.u.). These data clearly establish that the position of the magnetic peak is ${\bf Q}=(1,0,0.5)$, which corresponds to the propagation vector $\bf{k}$=(0, 0, 0.5). The magnetic ground state is thus antiferromagnetic (commensurate). The temperature dependence of the intensity at the peak position ${\bf Q}=(1,0,0.5)$ is shown in Fig. 8a. The signal vanishes at $T_{N2}$ = 23.5 K. In order to determine the magnetic structure, a collection of magnetic peaks was undertaken and is limited to eight peaks (7 independent ones) due to the weak signal associated with the small sample for neutron diffraction experiment. The scale factor and the extinction correction are obtained from nuclear peaks (44 independent reflections). Among the structural parameters given in Table 1, the isotropic thermal displacements (except for vanadium which has a very small coherent neutron scattering cross-section) and the $z$ position of Sn3 are refined yielding values consistent with the X-ray refinement. The magnetic form factor $f(\bf{Q})$ is fixed to the one of U3+ in the dipolar approximation $f(Q)$=$j_{0}$($Q$)+$C_{2}$$j_{2}$($Q$) using the value $C_{2}$=1.75 where $j_{0}(Q)$ and $j_{2}(Q)$ are spherical Bessel functions [41]. All the magnetic reflections with ${\bf Q}$ parallel to the $c$-axis, namely (0, 0, 0.5), (0, 0, 1.5) and (0, 0, 2.5), have zero intensity, which according to neutron scattering selection rules points out to the fact that the magnetic moment must lie along the $c$-axis. The magnitude of the moment is calculated to be 1.43(3) $\mu_{B}$. The refined magnetic structure is shown in Fig.8b.
@@
Figure 7: Momentum transfer cuts obtained on D23 at 1.41 K (full blue circles) and 30 K (open red circles) along (a) ${\bf Q}=(Q_{H},0,0.5)$, (b) along ${\bf Q}=(1,Q_{K},0.5)$ and (c) ${\bf Q}=(1,0,Q_{H})$. Full lines are fits to gaussian and dashed lines serve as guides to the eyes d) Neutron intensity obtained on the XtremeD detector bank at 24.9 K for the sample rotation angle achieving diffraction condition for the peak (1, 0, 0.45) (green spot).
@@
Above $T_{N2}$, the search for the ordering wave-vector was performed on XtremeD at ILL. A representative acquisition on the detector bank is shown in Fig.7d for a temperature of 24.9 K. It is obtained by rotating the single crystal sample with a step of 0.2 and an acquisition time of approximatively 10 minutes. A peak is observed at a position corresponding to the momentum transfer ${\bf Q}=(1,0,0.45)$. This incommensurate position is found to be temperature dependent and further measurements were performed on D23 to investigate this using the PG monochromator. The obtained temperature variation of the incommensurate position is shown in Fig.8c through the temperature dependence of the neutron intensity as a function of $Q_{L}$. Hence the AFM1 phase is incommensurate at zero magnetic field (the evolution of the propagation vector under magnetic field is not addressed in the present study). The temperature dependence of the $Q_{L}$-direction integrated intensity is shown in Fig. 8a. The intensity collapses at $T_{N1}$=30 K. There is a temperature range near $T_{N2}$, where both antiferromagnetic and incommensurate phases coexist, in agreement with a transition of the first-order. In addition, the mapping capabilities of XtremeD allow to confirm that only the antiferromagnetic propagation vector ${\bf Q}=(1,0,0.5)$ is present at 1.5 K.
@@
Figure 8: a) Temperature dependence of the intensity at ${\bf Q}=(1,0,0.5)$ (left scale) and of the integrated intensity at ${\bf Q}=(1,0,Q_{L})$ (right scale). b) Sketch of the antiferromagnetic ground state structure (color of atoms as in Fig. 1) c) Color-coded temperature dependence of the neutron intensity as a function of temperature and $Q_{L}$.
@@
Among rare earth based RV6Sn6 compounds, the magnetic structures were determined for R= Gd by resonant X-ray diffraction [42] and for R=Tb, Dy, Ho, Er by neutron diffraction [43]. In GdV6Sn6, two transitions are observed at 3.8 K and 5.2 K. In the high temperature phase, only an incommensurate propagation vector, ${\bf k}$=(0, 0, $q_{IC}$), is observed, with a small temperature dependence of $q_{IC}$ ranging between 0.46 and 0.47. Below the second transition, it coexists with a commensurate component ${\bf k}$=(0, 0, 0.5) down to the lowest measured temperature of 2 K. A lock-in at lower temperatures is suggested. The magnetic moments lie presumably predominantly in the basal plane of the structure. TbV6Sn6, DyV6Sn6 and HoV6Sn6 are ferromagnetic materials with transition temperatures at 4.3 K, 3 K and 2.4 K respectively and moments along the $c$-axis, except for R=Dy, where a 20 tilt towards the $a$-axis is found. ErV6Sn6 is antiferromagnetic below 0.6 K with the propagation vector ${\bf k}$=(0, 0, 0.5) and the magnetic moments along the $a$-axis in the hexagonal plane. In these heavy rare earth compounds, the magnetic moment lies in the range 6-9 $\mu_{B}$ (while it is no determined for GdV6Sn6).
@@
The magnetic structure of UV6Sn6 is distinct from each one reported for its rare-earth counterparts: it adopts the antiferromagnetic propagation vector common to Gd and Er compounds but the moment direction found in the Tb and Ho ones. One can also notice the highest transition temperature by almost one order of magnitude as compared to the rare earth systems. The magnetic moment for 1.42(3) $\mu_{B}$ is in the typical range of values found in uranium intermetallics. The ground state antiferromagnetic phase is made of ferromagnetic planes, with presumably a strong exchange interaction given the positive Curie-Weiss temperature and with an antiferromagnetic stacking along the $c$-axis involving likely a smaller interaction. The observed full lock-in transition from the incommensurate to the antiferromagnetic structure with decreasing temperature could be related to a highest transition temperature as compared to the Gd system, where the lock-in is partial. An interesting and open question concerns the arrangement of the spins in the supercell, of which determination goes beyond the present work. One can mention the simplest structure where all the spins in the $z$=0 and $z$=1/2 layers (see notations above) of the supercell would be ferromagnetic. In this case the global structure can be seen as composed of antiferromagnetic blocks made of the magnetic structure as described here in the HfFe6Ge6 cell with ”sliding defects” in the structure at the boundary between $z$=0 and $z$=1/2 layers. Another related global description of the magnetic structure can be made in the YCo6Ge6-model, having a halving of the $c$-axis parameter and half occupancy of uranium sites (see above), where the magnetic arrangement is of type ”+ + - -” with a propagation vector ${\bf k}$=(0, 0, 0.25).

### v3

@@ III.2 Magnetization
The magnetization data collected from a single crystal of UV6Sn6, with a field of 1 T applied parallel and perpendicular to the crystallographic $c$ axis, is presented in Fig. 3. The magnetization as a function of temperature reveals a strong magnetic anisotropy, with the $c$ axis being the easy magnetization axis (Fig. 3a). Along the $ab$-plane, the magnetization remains low and nearly constant across the entire temperature range. Notably, the magnetization parallel to the $c$ axis exhibits a clear maximum at $T_{N_{1}}\approx 29$ K. A second transition occurring at approximately 23 K is also revealed as a kink in the magnetization curve (inset Fig. 3a and Fig. 12 in Appendix E). The effective magnetic moment, derived from the Curie-Weiss fit which holds down to about 70 K (Fig. 3b) along the $c$ axis, is found to be $3.89(1)$$\mu_{B}$/U. This value is slightly higher than the effective moment of the U3+ or U4+ configurations. The polycrystalline averaged susceptibility ($\chi_{avg}=\frac{1}{3}\chi_{c}+\frac{2}{3}\chi_{ab}$) does not follow the Curie-Weiss behavior over a large temperature range, indicating the importance of the crystal electric field (CEF). A detailed CEF analysis was recently reported [33]. The results indicate that the first excited CEF level is between 18 to 25 meV, consistent with an isolated ground state of easy-axis moments [33].
The magnetization isotherms $M(H)$ collected at 2 K are shown in Fig. 3c. When the magnetic field is applied parallel to the $c$ axis, a clear hysteresis effect is observed, along with two notable changes in the magnetization. There is a sharp change in slope at approximately 3 T, followed by a tendency towards saturation starting around 4 T. The magnetization at 6 T is roughly 1 $\mu_{B}$, which is lower than the ordered moment of the free U ion. However, as discussed below, the magnetization at 6 T does not correspond to the fully polarized moment.
Figure 3: (a) Temperature dependence of the magnetization under an applied magnetic field of 1 T parallel and perpendicular to the $c$ axis. Inset: Enlarged view at low temperatures. The arrows indicate the anomalies at the two transitions. (b) Temperature dependence of the inverse susceptibility at 1 T along the c-axis and for a polycrystalline average: $\chi_{avg}=\frac{1}{3}\chi_{c}+\frac{2}{3}\chi_{ab}$. The dashed line is a fit to the Curie-Weiss law. (c) Magnetic field dependence of the magnetization at 2 K for the field parallel and perpendicular to the $c$ axis.
@@ III.3 Resistivity
The temperature dependence of the normalized electric resistivity at magnetic fields of 0 and 9 T applied along the $c$ axis is presented in Fig. 4. The resistivity ratio $RRR=R(300{\rm K})/R(1.8{\rm\,K})\approx 3$ is rather low. A similar value was reported recently [33]. On cooling the resistivity first increases slightly and shows a maximum around 180 K. At high temperatures, no magnetoresistance occurs and down to 90 K the temperature dependence at 9 T is similar to that at zero magnetic field. On further cooling, the resistivity shows a marked kink at the onset of the magnetic order near $T_{N1}\simeq 29~{}K$ (see Fig. 4b). At 6 T, the magnetic transition is still visible by a clear kink, while it disappears at 9 T.
Figure 4c presents the derivative of the normalized resistivity as a function of temperature. It shows a mean-field-like anomaly at the transition, as expected from the Fisher-Langer behavior $d\rho/dT\propto C_{P}$, often observed in metallic systems [41, 42]. In particular, we do not observe an increase of resistivity upon lowering the temperature through the magnetic transition, unlike other incommensurate antiferromagnets due to the opening of a superzone gap in the electronic band structure [43, 44, 45, 46, 47, 48]. This may be due to the electrical current direction being perpendicular to the magnetic wave vector (see neutron scattering and the magnetic structure discussion below).
@@
Figure 4: (a) Temperature dependence of the electrical resistivity at magnetic fields of 0 T and at 9 T applied along the $c$ axis, normalized to the value at 300 K. (b) Zoom of the low temperature range for fields of 0, 6, and 9 T. The arrows indicate the magnetic ordering temperature $T_{N1}$. (c) Temperature derivative of the resistivity at zero field as a function of temperature. The arrows show the magnetic transition temperatures $T_{N1}$ and $T_{N2}$.
@@ III.4 Heat capacity
Figure 5: Specific heat divided by temperature vs temperature of UV6Sn6 at zero magnetic field. The arrows indicate the position of $T_{N1}=28.8$ K and $T_{N2}=23$ K. The inset shows $C_{p}$ as a function of $T^{2}$. The dashed line is a linear fit to $C_{p}/T=\gamma+\beta T^{2}$ and we determine $\gamma=15~{}$mJmol-1K-2 and $\beta=2.95$ mJmol-1K-4.
Figure 6: (a) Specific heat divided by temperature at selected T values as a function of increasing magnetic fields. (b) Field dependence of the specific heat at 10 K in an enlarged view. (c) Magnetic phase diagram for field $H\parallel c$ from the field dependent specific heat measurements (increasing fields). (Lines are guides to the eyes.) An additional phase boundary within the AFM1 region was inferred from thermal expansion measurements [33].
@@
A combined plot of $dM/dT$, $C_{p}/T$, and $d\rho/dT$ as a function of $T$ is presented in Fig. 12 in Appendix E. $T_{N1}$ and $T_{N2}$ are confirmed from the three data sets.
In Fig. 6a, we show the field dependence of the specific heat at various temperatures for increasing magnetic fields. As shown explicitly in Fig. 6b for the curve at 10 K, a first increase of the specific heat is observed at 2.5 K while a very marked anomaly occurs at 7.7 K. While the low field anomaly for all temperatures is very small, the high field anomaly appears as a very pronounced peak with a first-order like behavior at 15 and 20 K. From these anomalies, a magnetic phase diagram can be proposed as illustrated in Fig. 6c. The low field anomaly is associated with a change in the magnetic structure, as discussed below.
@@ III.5 Neutron scattering and magnetic structure
Mapping of the reciprocal space along the high-symmetry lines of the Brillouin zone was carried out on D23 with the copper monochromator to identify the magnetic ordering wave vector at 1.41 K. The outcome is the identification of a single wave vector that is described below. Given the complexity of the UV6Sn6 crystal structure (Fig. 1), we approach the magnetic structure characterization from the HfFe6Ge6-type parent cell. Figure 7 shows scans around the momentum transfer ${\bf Q}=(1,0,0.5)$ along the directions $Q_{H}$ (panel a), $Q_{K}$ (panel b) and $Q_{L}$ (panel c) respectively, at 1.5 K and 30 K. The coordinates of the momentum transfer $\bf{Q}$ are given in reciprocal lattice units (r.l.u.). These data clearly establish that the position of the magnetic peak is ${\bf Q}=(1,0,0.5)$, which corresponds to the propagation vector $\bf{k}$=(0, 0, 0.5). The magnetic ground state is thus antiferromagnetic (commensurate). The same magnetic wave vector was recently reported from powder neutron diffraction at 2 K [33]. The temperature dependence of the intensity at the peak position ${\bf Q}=(1,0,0.5)$ is shown in Fig. 8a. The signal vanishes at $T_{N2}$ = 23.5 K. To determine the magnetic structure, a collection of magnetic peaks was undertaken and is limited to eight peaks (7 independent ones) due to the weak signal associated with the small sample for neutron diffraction experiment. The scale factor and the extinction correction are obtained from nuclear peaks (44 independent reflections). Among the structural parameters given in Table 1, the isotropic thermal displacements (except for vanadium which has a very small coherent neutron scattering cross section) and the $z$ position of Sn3 are refined yielding values consistent with the x-ray refinement. The magnetic form factor $f(\bf{Q})$ is fixed to the one of U3+ in the dipolar approximation $f(Q)$=$j_{0}$($Q$)+$C_{2}$$j_{2}$($Q$) using the value $C_{2}$=1.75, where $j_{0}(Q)$ and $j_{2}(Q)$ are spherical Bessel functions [50]. All the magnetic reflections with ${\bf Q}$ parallel to the $c$ axis, namely, (0, 0, 0.5), (0, 0, 1.5) and (0, 0, 2.5), have zero intensity, which, according to neutron scattering selection rules point to the fact that the magnetic moment must lie along the $c$ axis. The magnitude of the moment is calculated to be 1.43(3) $\mu_{B}$. The refined magnetic structure is shown in Fig.8b.
@@
Figure 7: Momentum transfer cuts obtained on D23 at 1.41 K (full blue circles) and 30 K (open red circles) along (a) ${\bf Q}=(Q_{H},0,0.5)$, (b) along ${\bf Q}=(1,Q_{K},0.5)$, and (c) ${\bf Q}=(1,0,Q_{H})$. Full lines are fits to Gaussian and dashed lines serve as guides to the eyes. (d) Neutron intensity obtained on the XtremeD detector bank at 24.9 K for the sample rotation angle achieving diffraction condition for the peak (1, 0, 0.45) (green spot).
@@
Above $T_{N2}$, the search for the ordering wave vector was performed on XtremeD at ILL. A representative acquisition on the detector bank is shown in Fig.7d for a temperature of 24.9 K. It is obtained by rotating the single crystal sample with a step of 0.2 and an acquisition time of approximatively 10 minutes. A peak is observed at a position corresponding to the momentum transfer ${\bf Q}=(1,0,0.45)$. This incommensurate position is found to be temperature dependent and further measurements were performed on D23 to investigate this using the PG monochromator. The obtained temperature variation of the incommensurate position is shown in Fig.8c through the temperature dependence of the neutron intensity as a function of $Q_{L}$. Hence, the AFM1 phase is incommensurate at zero magnetic field (the evolution of the propagation vector under magnetic field where additional phase boundaries are observed [33] is not addressed in the present paper). The temperature dependence of the $Q_{L}$-direction integrated intensity is shown in Fig. 8a. The intensity collapses at $T_{N1}\approx 29$ K. There is a temperature range near $T_{N2}$, where both antiferromagnetic and incommensurate phases coexist, in agreement with a transition of the first order. In addition, the mapping capabilities of XtremeD allow us to confirm that only the antiferromagnetic propagation vector ${\bf Q}=(1,0,0.5)$ is present at 1.5 K.
@@
Figure 8: (a) Temperature dependence of the intensity at ${\bf Q}=(1,0,0.5)$ (left scale) and of the integrated intensity at ${\bf Q}=(1,0,Q_{L})$ (right scale). (b) Sketch of the antiferromagnetic ground state structure (color of atoms as in Fig. 1). (c) Color-coded temperature dependence of the neutron intensity as a function of temperature and $Q_{L}$.
@@
Among rare-earth based RV6Sn6 compounds, the magnetic structures were determined for $R$ = Gd by resonant x-ray diffraction [51] and for $R$ = Tb, Dy, Ho, Er by neutron diffraction [52]. In GdV6Sn6, two transitions are observed at 3.8 K and 5.2 K. In the high temperature phase, only an incommensurate propagation vector, ${\bf k}$=(0, 0, $q_{IC}$), is observed, with a small temperature dependence of $q_{IC}$ ranging between 0.46 and 0.47. Below the second transition, it coexists with a commensurate component ${\bf k}$=(0, 0, 0.5) down to the lowest measured temperature of 2 K. A lock-in at lower temperatures is suggested. The magnetic moments lie presumably predominantly in the basal plane of the structure. TbV6Sn6, DyV6Sn6 and HoV6Sn6 are ferromagnetic materials with transition temperatures at 4.3 K, 3 K and 2.4 K respectively and moments along the $c$ axis, except for $R$ = Dy, where a 20 tilt toward the $a$ axis is found. ErV6Sn6 is antiferromagnetic below 0.6 K with the propagation vector ${\bf k}$=(0, 0, 0.5) and the magnetic moments along the $a$ axis in the hexagonal plane. In these heavy rare-earth compounds, the magnetic moment lies in the range 6-9 $\mu_{B}$ (while it is not determined for GdV6Sn6).
@@
The magnetic structure of UV6Sn6 is distinct from each one reported for its rare-earth counterparts: it adopts the antiferromagnetic propagation vector common to Gd and Er compounds but the moment direction found in the Tb and Ho ones. One can also notice the highest transition temperature by almost one order of magnitude as compared to the rare-earth systems. The magnetic moment for 1.42(3) $\mu_{B}$ is in the typical range of values found in uranium intermetallics. The ground-state antiferromagnetic phase is made of ferromagnetic planes, with presumably a strong exchange interaction given the positive Curie-Weiss temperature and with an antiferromagnetic stacking along the $c$ axis involving likely a smaller interaction. The observed full lock-in transition from the incommensurate to the antiferromagnetic structure with decreasing temperature could be related to the highest transition temperature as compared to the Gd system, where the lock-in is partial. An interesting and open question concerns the arrangement of the spins in the supercell, of which determination goes beyond the present paper. One can mention the simplest structure where all spins in the $z$ = 0 and $z$ = 1/2 layers (see notations above) of the supercell would be ferromagnetic. In this case, the global structure can be seen as composed of antiferromagnetic blocks made of the magnetic structure as described here in the HfFe6Ge6 cell with ”sliding defects” in the structure at the boundary between $z$ = 0 and $z$ = 1/2 layers. Another related global description of the magnetic structure can be made in the YCo6Ge6 model, having a halving of the $c$ axis parameter and half occupancy of uranium sites (see above), where the magnetic arrangement is of type ”+ + - -” with a propagation vector ${\bf k}$ = (0, 0, 0.25).